# Pyramid-Monozone Synergistic Grasping Policy in Dense Clutter

### v1

@@ Abstract
Grasping a diverse range of novel objects from dense clutter poses a great challenge to robots because of the occlusion among these objects. In this work, we propose the Pyramid-Monozone Synergistic Grasping Policy (PMSGP) that enables robots to cleverly avoid most occlusions during grasping. Specifically, we initially construct the Pyramid Sequencing Policy (PSP) to sequence each object in the scene into a pyramid structure. By isolating objects layer-by-layer, the grasp candidates will focus on a single layer during each grasp. Then, we devise the Monozone Sampling Policy (MSP) to sample the grasp candidates in the top layer. Through this manner, each grasp will target the topmost object, thereby effectively avoiding most occlusions. We perform more than 7000 real-world grasping among 300 novel objects in dense clutter scenes, demonstrating that PMSGP significantly outperforms seven competitive grasping methods. All grasping videos are available at: https://www.youtube.com/@chenghaoli4532/playlists.
@@ I INTRODUCTION
Grasping a wide range of novel objects reliably can benefit applications in warehousing, manufacturing, retail, and services. However, when these novel objects are densely cluttered, occlusions make this task extremely challenging. As shown in Fig. 1, the object shape incompleteness caused by occlusions can greatly undermine grasping reliability.
One approach to dense clutter grasping is to build a database of grasps on three-dimensional (3D) object models, using performance metrics based on geometry and physics [1, 2], along with stochastic sampling to model uncertainty [3]. This analytical method relies on a perception system to register sensor data with known objects, but it often fails to generalize to various novel objects in dense clutter. To solve this problem, another approach [4, 5] uses machine learning to train function approximators, such as deep neural networks (DNN), to predict the success probability (quality score-based sampling) of grasp candidates from images by leveraging large datasets of empirical successes and failures. However, quality score-based sampling almost does not consider the occlusion problem, which potentially leads to the optimal grasp being predicted on occluded objects prone to collision.
@@
A common solution is to first segment all objects in the scene to create a mask, and then use this mask to guide the sampling of the quality score [6, 7]. This involves evaluating the relation of each instance (object) and determining whether their grasp candidates might result in collisions. But, this method is constrained by the quality score itself, that is, the quality score does not guarantee that every instance has a viable grasp candidate, even if the instance is not occluded. Furthermore, this method typically requires classifying and segmenting all objects in the scene, and these segmented objects may also have incomplete shapes due to occlusion. Thus, this method still does not effectively solve the occlusion problem and essentially trades high costs for limited improvements.
Figure 1: Dense clutter grasping example: Object occlusions are anywhere in this scene (highlighted in green). The objects are often partially visible and closely piled, so grasping them directly can easily fail.
@@
Regarding these problems, we consider that since the robot typically grasps one object at a time, why not identify an unoccluded region before each grasping and only sample within this region, which would not only be accurate but also avoid redundant analysis?
@@
In this paper, we present a novel grasping policy, called the Pyramid-Monozone Synergistic Grasping Policy (PMSGP), which first leverages the Pyramid Sequencing Policy (PSP) to sequence each object in the scene into a pyramid structure, meaning that before each grasp, only the top layer instance is segmented, thus the grasp candidates will focus on this layer. Then, through the Monozone Sampling Policy (MSP), the sampling will be also concentrated on this top-layer instance, more precisely, we do not use the mask to guide the quality score-based sampling, but directly sample these grasp candidates within the mask. This pyramid and monozone operation helps the robot cleverly avoid most occlusions during grasping, and as the grasping progresses, the extent of occlusion gradually decreases, thereby reducing the difficulty of subsequent grasps. Our contributions are as follows:
@@
    We simplify the complex occlusion problem in dense clutter grasping through first hierarchical, then locally processing ideas.
@@
    We propose a novel and cost-effective grasping policy based on pyramid sequencing and monozone sampling.
@@
    We demonstrate that our method far outperforms seven competitive methods among 300 novel objects in various dense clutter scenes. More importantly, even in extremely cluttered scenes with up to 100 novel objects, our method still exhibits high reliability.
@@ II Related Work
While many grasping frameworks exist, this work only focuses on vision-guided 4-DOF grasping with a parallel-jaw gripper. The 4-DOF grasp framework typically performs grasping in a top-down manner, where the robot moves along the $X$, $Y$ and $Z$-axes and rotates only around the $Z$-axis. During grasping, the parallel-jaw gripper will adjust to an appropriate width based on the size of the object perceived by the depth camera. It is mainly divided into analysis-based methods and learning-based methods, as follows.
II-A Analytic-based Grasping Methods
Analytical methods employ mathematical and physical models of geometry, kinematics, and dynamics to compute stable grasps [1, 2, 3]. However, these methods often fail to transfer to a diverse range of novel objects in dense clutter.
@@ II-B Learning-based Grasping Methods
Learning-based methods can generalize to various novel objects in dense clutter, which typically involve training a function approximator, such as DNN, to predict the success probability of grasp candidates from images by leveraging large datasets of empirical successes and failures. As a result, datasets play a crucial role in these methods. One human-labeled dataset is the Cornell Grasping Dataset [8], which contains around 1k RGB-D images and has been widely used to train grasping models based on convolutional neural networks (CNN) [9], such as [10, 11, 12, 13, 14]. However, this dataset is quite small and consists only of single-object images, which limits the dense clutter grasping capabilities of models trained on it.
@@
The Dex-Net series [4, 15, 16, 17, 18] made significant advancements by generating large synthetic datasets that incorporate various dense clutter scenes. Despite these advancements, this approach did not fully resolve the sim-to-real problem. GraspNet [5, 19, 20], in contrast, constructed a real-world dataset featuring 1 billion grasp labels and nearly 100k images with 190 different dense clutter scenes, supporting both 4-DOF and 6D grasping. This dataset enabled remarkable real grasping performance in dense clutter. However, the sampling used in the above methods relies on predicting the success probability (quality score) of the grasp candidates, which almost does not consider the occlusion problem, potentially leading to the optimal grasp being predicted on occluded objects prone to collision.
@@
Recently, [6, 7] proposed to segment all objects in a scene to create a mask and then to use this mask to guide the sampling of quality scores. This process involves evaluating the relation of each object and assessing whether its grasp candidates might result in collisions. However, the effectiveness of this method is also limited by the quality score itself. Specifically, the quality score does not ensure that every object has a viable grasp candidate, even when the object is not occluded. Additionally, this method typically requires classifying and segmenting all objects in the scene, and these segmented objects may still have incomplete shapes due to occlusion. Therefore, this method also fails to effectively address the occlusion issue and essentially exchanges high costs for only marginal improvements. Unlike this method, we first organize each object in the scene into a pyramid structure, segmenting only the topmost object during each grasp. By segmenting layers by layers, most grasp candidates can be restricted to a single layer. We then directly sample grasp candidates for the object in the top layer for each grasp. This allows the robot to cleverly avoid most occlusions during grasping.
@@ III Grasp Configuration
We have previously provided a brief introduction to the 4-DOF grasping framework. Now, we will elaborate on how the grasp configuration is represented in the image coordinate system and its conversion to the robot end effector coordinate system (eye-in-hand grasping). Specifically, we adopt the same grasp configuration in [11], which is composed of parameters $(x,y,w,h,\theta)$ forming a rotated box. Here, $(x,y)$ represents the center of the box, $w$ and $h$ denote the width and height of the box, and $\theta$ represents the angle of the box relative to the horizontal direction.
Since $h$ is used only for visual representation and not in the conversion process, we denote the grasp configurations in the image and robot end effector coordinate systems as $g_{i}(x,y,w,\theta)$ and $g_{r}(x_{r},y_{r},w_{r},\theta_{r})$, respectively. Here, $(x_{r},y_{r},z_{r})$ represents the grasp position in the robot end effector coordinate system, $w_{r}$ is the opening width of the parallel jaw gripper, and $\theta_{r}$ is the rotation angle of the gripper relative to the $Z$ axis. The parameter conversion between $g_{i}$ and $g_{r}$ is described by the following equations:
@@
Figure 2: The pipline of PMSGP: First, apply Top View Alignment (TVA) to align the depth camera center above the topmost object, and use Cross-prompted Segmentation (CPS) to isolate the object from the scene, obtaining ${M_{r}}$. Next, input both ${M_{r}}$ and the depth image into the grasping model to generate grasp candidates $G$, refining them with Grasp Angle Calibration (GAC) to produce $G^{\prime}$. Finally, sample the optimal grasp $g^{*}$ using Monozone Grasp Analysis (MGA), and further optimize it into $g^{*}_{f}$ as the final grasp.
The conversion process is divided into two parts. The first part involves converting $(x,y)$: using depth information ($d$) and the camera’s intrinsic parameters ($f_{x}$, $f_{y}$ for focal lengths and $c_{x}$, $c_{y}$ for the image center coordinates), we convert $(x,y)$ from the image coordinate system to the camera coordinate system $(x_{c},y_{c},z_{c})$. This is followed by converting $(x_{c},y_{c},z_{c})$ to the robot end effector coordinate system $(x_{r},y_{r},z_{r})$ using the positional transformation relation $T_{rc}$. The second part involves converting $(w,\theta)$ into $(w_{r},\theta_{r})$ using the projection function $P_{ri}$.
The intrinsic parameters and depth information are directly obtained from the depth camera, and $T_{rc}$ is derived from offline eye-in-hand calibration. $T_{rc}$ depends on the relative orientations (excluding rotation) of the $X$, $Y$, and $Z$ axes between the camera coordinate system and the robot end effector coordinate system. Finally, the projection function $P_{ri}$ allows for manual adjustment of the linear relationship between the gripper’s width $w_{r}$ and rotation $\theta_{r}$ relative to the grasp box’s width $w$ and rotation $\theta$.
Once the final grasp pose in the robot end effector coordinate system $(x_{r},y_{r},z_{r},\theta_{r},\theta_{x}^{*},\theta_{y}^{*})$ is obtained, where $\theta_{x}^{*}$ and $\theta_{y}^{*}$ represent the constant rotations relative to the $X$-axis and the $Y$-axis, the gripper will move to the target pose using inverse kinematics and open up to the width $w_{r}$ to grasp.
@@ IV Proposed Method
We propose a novel grasping policy, the Pyramid-Monozone Synergistic Grasping Policy (PMSGP), designed to improve grasping reliability in dense clutter, As illustrated in Fig. 2. PMSGP is composed of two main modules: the Pyramid Sequencing Policy (PSP) and the Monozone Sampling Policy (MSP). The PSP is used for layering objects in cluttered scenes, with submodules for Top View Alignment (TVA) and Cross-prompted Segmentation (CPS). The MSP samples the topmost objects based on the layering results from the PSP, which includes Grasp Angle Calibration (GAC), Monozone Grasp Analysis (MGA), and Optimal Grasp Refinement (OGR). Overall, The MSP is inherently dependent on the PSP, and the submodules within both are linked. Next, we will provide a detailed explanation for them.
@@ IV-A Pyramid Sequencing Policy (PSP)
How to determine an appropriate grasp sequence in dense clutter to decrease the difficulty of grasping? In this context, we introduce the Pyramid Sequencing Policy (PSP), a method that organizes objects within a dense cluttered scene into a pyramid structure, where each layer contains a single object, and the grasp candidate will focus only on the object at the highest layer each time. It is analogous to disassembling a stack of blocks, but in reverse order; rather than building from the ground up, we sequentially remove objects from the top down. This ensures stability and prevents collapse during the disassembly process.

### v2

@@ Abstract
Grasping a diverse range of novel objects in dense clutter poses a great challenge to robotic automation mainly due to the occlusion problem. In this work, we propose the Pyramid-Monozone Synergistic Grasping Policy (PMSGP) that enables robots to effectively handle occlusions during grasping. Specifically, we initially construct the Pyramid Sequencing Policy (PSP) to sequence each object in cluttered scenes into a pyramid structure. By isolating objects layer-by-layer, the grasp detection model is allowed to focus on a single layer during each grasp. Then, we devise the Monozone Sampling Policy (MSP) to sample the grasp candidates in the top layer. Through this manner, each grasp targets the topmost object, thereby effectively avoiding most occlusions. We performed more than 7,000 real-world grasping in densely cluttered scenes with 300 novel objects, demonstrating that PMSGP significantly outperforms seven competitive grasping methods. More importantly, we tested the grasping performance of PMSGP in extremely cluttered scenes involving 100 different household goods, and found that PMSGP pushed the grasp success rate to 84.9%. To the best of our knowledge, no previous work has demonstrated similar performance. All grasping videos are available at: https://www.youtube.com/@chenghaoli4532/playlists.
Index Terms:
Robot grasping, grasp detection, deep learning, dense clutter, grasp sequence policy, grasp sampling.
@@ I INTRODUCTION
Grasping a wide range of novel objects can benefit applications such as warehousing, manufacturing, retail, and services. However, when novel objects are densely cluttered, occlusions can make this task extremely challenging. As shown in Fig. 1, incomplete object shapes due to occlusions may greatly undermine grasping capability.
One approach to dense clutter grasping is to build a database of grasps on three-dimensional (3D) object models, using performance metrics based on geometry and physics [1, 2], along with stochastic sampling for uncertainty quantification [3]. Although these techniques provide robust solutions in structured environments, they are fundamentally constrained by their dependence on complete 3D object models. In real-world applications, such models are not always accessible, especially when robots operate in unstructured environments with numerous unknown objects. Additionally, building accurate 3D models is often labor-intensive, expensive, and time-consuming, demanding significant manual effort for each unique object. These limitations underscore the need for more adaptable and efficient approaches to robot grasping that can cope with uncertainty and variations in object geometry. In response to these challenges, there has been a new approach [4, 5] leveraging deep neural networks (DNN) [6, 7, 8, 9, 10] to train the function approximator. The approximator predicts the best grasp candidate from images based on quality score-based sampling utilizing large datasets of empirical successes and failures. This approach can generalize to novel objects at a significantly low cost. However, the quality score-based sampling used in these methods rarely considers the occlusion problem, which potentially leads the optimal grasp predicted on occluded objects to a collision.
@@
A common solution to the problem mentioned above is to first segment all objects in the scene to create a mask, and then use this mask to guide the sampling of the quality score [11, 12]. This involves evaluating the relation of each instance (object) and determining whether their grasp candidates might result in collisions. But, this method is constrained by the quality score itself, that is, the quality score does not guarantee that every instance has a viable grasp candidate, even though the instance is not occluded. Furthermore, this method typically requires classifying and segmenting all objects in the scene, and these segmented objects may also have incomplete shapes due to occlusion. Thus, this method still does not effectively solve the occlusion problem and essentially trades high costs for limited improvements.
Figure 1: Grasping in dense clutter: Objects are occluded to one another (highlighted in green lines) and partially visible, which can easily lead to failed grasping.
@@
Now we look at the grasp problem in clutter from a different perspective. Since the robot typically grasps one object at a time, why not identify an unoccluded region before each grasp attempt and only sample within this region, which would not only be accurate but also avoid redundant analysis?
@@
In this paper, we present a novel grasping policy, called the Pyramid-Monozone Synergistic Grasping Policy (PMSGP), which leverages the Pyramid Sequencing Policy (PSP) to sequence each object in the scene into a pyramid structure, meaning that before each grasp, only the top layer instance is segmented. This allows the grasp detection model to focus on the top layer. Then, through the Monozone Sampling Policy (MSP), the grasp sampling is also concentrated on the top-layer instance. More precisely, we do not use the mask to guide the quality score-based sampling, but directly sample these grasp candidates within the mask. This pyramid and monozone operation help the robot avoid most occlusions during grasping, and as the grasping progresses, the extent of occlusion gradually decreases, thereby reducing the difficulty of subsequent grasps. Our contributions are as follows:
@@
    We propose a novel dense clutter grasping policy incorporating pyramid sequencing and monozone sampling policies.
@@
    We restructure the problem of grasping novel objects in dense clutter into a hierarchically layered grasp detection problem, each layer of which contains only a single object. This allows sampling grasp candidates from each single layer.
@@
    We conduct extensive real-world grasping experiments, and demonstrate that our method far outperforms seven competitive methods among 300 novel objects in various densely cluttered scenes. More importantly, even in extremely cluttered scenes with up to 100 household goods, our method still exhibits a high grasp success rate, reaching 84.9%.
This paper is organized into the following sections. Section II (Related Work) provides a review of traditional grasping methods and learning-based grasping methods. Section III (Grasp Configuration) describes the 4-DOF grasp configuration, and how to transform it from the image coordinates to the robot end effector coordinates. Section IV (Proposed Method) provides an overview of PMSGP and makes a detailed description of its two components (PSP and MSP), as well as each submodule of each component. Section V (Experiments) shows the real grasping results of our method and two groups of baseline methods in dense clutter and performed a failure analysis. Finally, Section VI (Conclusion) summarizes the work of this paper and provides prospects for future research.
@@ II Related Work
While many grasping frameworks exist, this work only focuses on vision-guided 4-DOF grasping with a parallel-jaw gripper. The 4-DOF grasp framework typically performs grasping in a top-down manner, where the robot moves along the $X$, $Y$, and $Z$-axis and rotates only around the $Z$-axis. During grasping, the parallel-jaw gripper will adjust its opening stroke based on the size of the object perceived by the depth camera. It is mainly divided into traditional methods and learning-based methods as follows.
II-A Traditional Grasping Methods
Traditional grasping methods rely on mathematical and physical models that describe the geometry, kinematics, and dynamics of objects [1, 2, 3]. These methods typically assume access to a detailed 3D model of the object, which is used to compute stable grasps. For example, Gallegos et al. [13] optimized grasp strategies by leveraging both a known 3D model of the object and predefined contact points for the robot gripper. Similarly, Pokorny et al. [14] proposed grasping spaces, where objects could be mapped to these spaces to identify suitable grasps. While these techniques offer robust solutions in controllable structured environments, they are inherently limited by their reliance on complete 3D object models. In real-world scenarios, such models may not always be available, particularly when robots are deployed in uncontrollable unstructured environments with many unknown objects. Moreover, the process of constructing accurate 3D models is often labor-intensive, costly, and time-consuming, requiring extensive manual effort for each unique object. Therefore, these constraints highlight the need for more adaptable and efficient approaches to robot grasping that can handle uncertainty and variability in object geometry.
@@ II-B Learning-based Grasping Methods
Learning-based methods can generalize to various novel objects in dense clutter, which typically involve training a function approximator, such as DNN, to predict the success probability of grasp candidates from images by leveraging large datasets of empirical successes and failures. For that reason, datasets play a crucial role in these methods. One human-labeled dataset is the Cornell Grasping Dataset [15], which contains around 1,000 RGB-D images and has been widely used to train grasping models, such as [17, 18, 19, 20, 21, 22, 23], based on convolutional neural networks (CNN) [16]. However, this dataset is quite small and consists only of single-object images, which limits the dense clutter grasping capabilities of models trained on it.
@@
The Dex-Net series [4, 24, 25, 26, 27] made significant advancements by generating large synthetic datasets that incorporate various dense clutter scenes. Despite these advancements, this approach did not fully resolve the sim-to-real problem. GraspNet [5, 28, 29], in contrast, constructed a real-world dataset featuring one billion grasp labels and nearly 100,000 images with 190 different dense clutter scenes, supporting both 4-DOF and 6-DOF grasping. This dataset enabled remarkable real grasping performance in dense clutter. However, the sampling used in the above methods relies on predicting the success probability (quality score) of the grasp candidates, which almost does not consider the occlusion problem, potentially leading the optimal grasp predicted on occluded objects to a collision.
@@
Recently, several works proposed to segment all objects in a scene to create a mask that can guide the sampling of quality scores [11, 12]. These works involve evaluating the relationships between objects and assessing whether each grasp candidate might result in collisions. However, the effectiveness of this approach is also limited by the quality score itself. Specifically, the quality score does not ensure that every object has a viable grasp candidate, even when the object is not occluded. Additionally, it typically requires classifying and segmenting all objects in the scene, and these segmented objects may still have incomplete shapes due to occlusions. Therefore, this approach also fails to effectively address the occlusion issue and essentially exchanges high costs for only marginal improvements. Unlike the above works, we first organize each object in the scene into a pyramid structure, segmenting only the topmost object during each grasp. Segmenting in a layer-by-layer manner, most grasp candidates can be restricted to a single layer. We then sample grasp candidates for the object in the top layer for each grasp. This allows the robot to avoid most occlusions during grasping.
Figure 2: Pipeline of PMSGP: Firstly, conducting Top View Alignment (TVA) to center the initial view $V$ of depth camera on the topmost object to get view $V^{\prime\prime\prime}$, and segment this topmost object by the center $V^{\prime\prime\prime}(c)$ (green point) of this view as prompt to obtain initial segmented RGB image (emphasized with green lines) with mask $M_{f}$. Then, calculate two pairs of most distant points ($p_{m}$ (red point), $p_{m}^{\prime}$ (red point), ${p_{m_{p}}}$ (blue point), and ${p_{m_{p}}^{\prime}}$ (blue point)) based on the edge of $M_{f}$, and using these points to make Cross-prompted Segmentation (CPS) to optimize $M_{f}$ to get ${M_{r}}$. In step three, the segmented RGB image with mask ${M_{r}}$ and the depth image within view $V^{\prime\prime\prime}$ are fed into the grasp detection (GD) model to generate initial grasp candidates $G$, followed by Grasp Angle Calibration (GAC) to obtain calibrated grasp candidates $G^{\prime}$. After GAC, $G^{\prime}$ will be through Monozone Grasp Analysis (MGA) to find the optimal grasp $g^{*}$. Finally, $g^{*}$ is optimized by Optimal Grasp Refinement (OGR) to transfer it to the final grasp $g^{*}_{f}$.
@@ III Grasp Configuration
Now, we elaborate on how the 4-DOF grasp configuration is represented in the image coordinate system and its conversion to the robot end effector coordinate system (eye-in-hand grasping). Specifically, we adopt the same grasp configuration in [18], which is composed of parameters $(x,y,w,h,\theta)$ forming a rotated box. Here, $(x,y)$ represents the center of the box, $w$ and $h$ denote the width and height of the box, and $\theta$ represents the angle of the box relative to the horizontal direction.
Since $h$ is used only for visual representation and not in the conversion process, we denote the grasp configurations in the image and robot end effector coordinate systems as $g_{i}(x,y,w,\theta)$ and $g_{r}(x_{r},y_{r},z_{r},w_{r},\theta_{r})$, respectively. Here, $(x_{r},y_{r},z_{r})$ represents the grasp position in the robot end effector coordinate system, $w_{r}$ is the opening stroke of the parallel jaw gripper, and $\theta_{r}$ is the rotation angle of the gripper relative to the $Z$ axis. The parameter conversion between $g_{i}$ and $g_{r}$ is described in Eq. 1, Eq. 2, and Eq. 3.
The conversion process is divided into two parts. The first part involves converting $(x,y)$: using depth information ($d$) and the camera’s intrinsic parameters ($f_{x}$, $f_{y}$ for focal lengths and $c_{x}$, $c_{y}$ for the image center coordinates), we convert $(x,y)$ from the image coordinate system to the camera coordinate system $(x_{c},y_{c},z_{c})$. This is followed by converting $(x_{c},y_{c},z_{c})$ to the robot end effector coordinate system $(x_{r},y_{r},z_{r})$ using the positional transformation relationship $T_{rc}$. The second part involves converting $(w,\theta)$ into $(w_{r},\theta_{r})$ using the projection function $P_{ri}$.
@@
The intrinsic parameters and depth information are directly obtained from the depth camera, and $T_{rc}$ is derived from offline eye-in-hand calibration. $T_{rc}$ depends on the relative orientations (excluding rotation) of the $X$, $Y$, and $Z$-axis between the camera coordinate system and the robot end effector coordinate system. Finally, the projection function $P_{ri}$ allows for manual adjustment of the linear relationship between the gripper stroke $w_{r}$ and rotation $\theta_{r}$ relative to the grasp box’s width $w$ and rotation $\theta$.
Once the final grasp pose in the robot end effector coordinate system $(x_{r},y_{r},z_{r},\theta_{r},\theta_{x}^{*},\theta_{y}^{*})$ is obtained, where $\theta_{x}^{*}$ and $\theta_{y}^{*}$ represent the constant rotations relative to the $X$-axis and the $Y$-axis, the gripper is moved to the target pose using inverse kinematics and its stroke is kept to the width $w_{r}$.
@@ IV Proposed Method
We propose a novel grasping policy, the Pyramid-Monozone Synergistic Grasping Policy (PMSGP), designed to improve the grasping performance of parallel-jaw grippers in dense clutter as illustrated in Fig. 2. PMSGP is composed of two main modules: the Pyramid Sequencing Policy (PSP) and the Monozone Sampling Policy (MSP). The PSP is used for layering objects in cluttered scenes, with submodules for Top View Alignment (TVA) and Cross-prompted Segmentation (CPS). The MSP samples the topmost objects based on the layering results from the PSP, which includes Grasp Angle Calibration (GAC), Monozone Grasp Analysis (MGA), and Optimal Grasp Refinement (OGR). Overall, the MSP is inherently dependent on the PSP, and the submodules within both are linked to each other. A detailed explanation of each module is given below.
@@ IV-A Pyramid Sequencing Policy (PSP)
It is important to know how to determine an appropriate grasp sequence in dense clutter to decrease the difficulty of grasping. In this context, we introduce the Pyramid Sequencing Policy (PSP), organizing objects in a cluttered scene into a pyramid structure, each layer of which contains only a single object, allowing the grasp detection model to focus on the object at the top layer each time. This is similar to reversing a sequence of piling up building blocks, that is, pulling out blocks from the top layer.